OpenStudy (anonymous):

Solve the following equation analytically. 2 log9(x) = log9(2) + log9(x + 40) (the 9's are base 9)

OpenStudy (anonymous):

\[2\log_9(x)=\log_9(2)+\log_9(x+40)\]

OpenStudy (anonymous):

\[2\log_{9}(x) = \log_{9}(2) + \log_{9}(x+40)\] \[\log_{9}(x^{2}) = \log_{9}(2(x+40))\] \[\log_{9}(x^{2}) = \log_{9}(2x+80)\] x^2-2x-80 (x-10)(x+8) x=10 or -8

OpenStudy (anonymous):

\[\log_9(x^2)=\log_9(2x+80)\] \[x^2=2x+80\] \[x^2-2x-80=0\] \[(x-10)(x+8)=0\] so \[x=10\] only because you cannot take the log of a negative number. so discard -8

OpenStudy (anonymous):

Actually -8 is extraneous because you can't have a negative log

OpenStudy (anonymous):

what edge said